A Certain rocket, initially at rest, is shot straight up with an acceleration of 6t m/s2 during the first seconds after blast-off, after which the engine cuts out and is subject


only to gravitational acceleration. How high will the rocket go? Use g=10 m/s2

1
Expert's answer
2021-12-16T00:03:01-0500

"V = u+at = 0+6t^2"


"V^2-u^2 = 2as"

"s = \frac{6t^2}{12t} =0.5t m"


The high will the rocket go = 0.5t meter.